ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

8. Analysis

1991 MS Classification: 30 Pérez Alvarez, Javier, and Fernández Arias, Arturo, Universidad Nacional de Educación a Distancia, Madrid, Spain Harmonic differentials on Klein Surfaces It is classical that on a Riemann surface of genus g, the vector spaces of real or complex harmonic differentials have dimension 2g with respect to the corresponding field. In this paper it is shown that for a nonclassical Klein surface of topological genus g and algebraic genus g̃, the real and complex harmonic differentials have real dimension g̃ and 2g̃ respectively.